# CADmium: Fine-Tuning Code Language Models for Text-Driven Sequential CAD Design

### v1

@@ 4 Experiments
This section provides an overview of our experimental setup and presents the evaluation of the proposed CAD generation pipeline. We conduct experiments across four datasets: the original Text2CAD dataset [2], our re-annotated DeepCAD dataset (CADmium), the CADPrompt dataset [11] containing 200 expert-annotated objects from DeepCAD [1], and the Fusion360 Reconstruction dataset [10] annotated using our pipeline.
@@
Additionally, for the Chamfer Distance, we observed that rare outliers with extremely high values skewed the mean, rendering it less informative. We therefore report the median CD as a more robust alternative. All metrics are accompanied by their standard deviation computed over individual test samples for a single model run.
@@ 4.1 Generalization to Human-Annotated Prompts (CADPrompt)
To assess our models’ ability to generalize to prompts authored by humans, we evaluated them on the CADPrompt dataset [11]. This dataset comprises expert-written natural language instructions for 200 DeepCAD samples, serving as a crucial test for out-of-distribution generalization. Table 1 presents the results, comparing our Qwen Coder 14B model trained on CADmium data against the original Text2CAD model as a key baseline. A corresponding common subset analysis is provided in Appendix B.
@@
Table 1: Performance comparison on the human-annotated CADPrompt dataset. Metrics are calculated on the set of valid meshes generated by each model individually.
Model Text2CAD Qwen Coder 14B
@@
Line F1 $\uparrow$ 0.67 $\pm$ 0.28 0.69 $\pm$ 0.31
Arc F1 $\uparrow$ 0.00 $\pm$ 0.00 0.22 $\pm$ 0.33
Circle F1 $\uparrow$ 0.31 $\pm$ 0.39 0.61 $\pm$ 0.43
Extrusion F1 $\uparrow$ 0.84 $\pm$ 0.20 0.89 $\pm$ 0.16
CD median $\downarrow$ 127.70 116.75
SIR $\downarrow$ 0.02 $\pm$ 0.13 0.11 $\pm$ 0.27
DangEL $\downarrow$ 0.20 $\pm$ 1.47 2.98 $\pm$ 10.01
SegE $\downarrow$ 0.09 $\pm$ 0.49 0.78 $\pm$ 2.25
FluxEE (×$\times 10^{2}$) $\downarrow$ 0.04 $\pm$ 0.43 2.24 $\pm$ 9.18
Watertightness (%) $\uparrow$ 95.74 83.80
@@
EECM $\uparrow$ 0.66 0.68
DMCD (×$\times 10^{3}$) $\downarrow$ 13.39 $\pm$ 12.14 8.42 $\pm$ 11.01
SD mean (×$\times 10^{2}$) $\downarrow$ 14.78 $\pm$ 13.98 7.02 $\pm$ 12.46
SD median (×$\times 10^{2}$) $\downarrow$ 11.15 1.54
@@ 4.2 Impact of Model Scale
A primary benefit of increasing model size is the consistent reduction in the Invalidity Ratio (IR). On the CADmium test set, IR drops progressively with larger models, and the 14B model achieves a notably lower IR on the more challenging Fusion360 dataset as well. Accuracy in reconstructing geometric primitives, as measured by F1 scores, generally improves or remains robust with increasing model size, particularly when compared on a common subset of generated objects (see Appendix B). Metrics related to fine-grained mesh quality, such as DangEL and SD, also tend to show improvements with larger models, suggesting better adherence to manifold properties.
Table 2: Impact of Qwen Coder model scale (1.5B, 3B, 7B, 14B parameters), trained on CADmium annotations, evaluated on the CADmium test set. Metrics are calculated on the set of valid meshes generated by each model individually for each dataset.
@@
Line F1 $\uparrow$ 0.90 $\pm$ 0.20 0.91 $\pm$ 0.19 0.91 $\pm$ 0.20 0.92 $\pm$ 0.18
Arc F1 $\uparrow$ 0.70 $\pm$ 0.38 0.74 $\pm$ 0.36 0.73 $\pm$ 0.37 0.78 $\pm$ 0.34
Circle F1 $\uparrow$ 0.87 $\pm$ 0.29 0.91 $\pm$ 0.23 0.89 $\pm$ 0.25 0.92 $\pm$ 0.22
Extrusion F1 $\uparrow$ 0.98 $\pm$ 0.07 0.98 $\pm$ 0.07 0.98 $\pm$ 0.06 0.99 $\pm$ 0.06
CD median $\downarrow$ 0.31 0.24 0.27 0.23
SIR $\downarrow$ 0.04 $\pm$ 0.16 0.05 $\pm$ 0.18 0.04 $\pm$ 0.16 0.04 $\pm$ 0.16
DangEL $\downarrow$ 1.18 $\pm$ 6.71 1.35 $\pm$ 6.59 1.10 $\pm$ 5.56 1.13 $\pm$ 6.03
SegE $\downarrow$ 0.48 $\pm$ 2.90 0.52 $\pm$ 2.35 0.48 $\pm$ 2.14 0.54 $\pm$ 5.03
FluxEE (×$\times 10^{2}$) $\downarrow$ 0.38 $\pm$ 3.61 0.49 $\pm$ 4.65 0.36 $\pm$ 3.32 0.39 $\pm$ 3.81
Watertightness (%) $\uparrow$ 90.38 90.31 90.54 90.78
@@
EECM $\uparrow$ 0.87 0.89 0.88 0.89
DMCD (×$\times 10^{3}$) $\downarrow$ 2.97 $\pm$ 7.80 2.68 $\pm$ 7.83 2.73 $\pm$ 8.08 2.67 $\pm$ 7.96
SD mean (×$\times 10^{2}$) $\downarrow$ 2.90 $\pm$ 8.43 2.19 $\pm$ 7.15 2.23 $\pm$ 7.30 1.99 $\pm$ 6.70
SD median (×$\times 10^{2}$) $\downarrow$ 0.00 0.00 0.00 0.00
Table 3: Impact of Qwen Coder model scale (1.5B, 3B, 7B, 14B parameters), trained on CADmium annotations, evaluated on the Fusion360 test set. Metrics are calculated on the set of valid meshes generated by each model individually for each dataset.
@@
Line F1 $\uparrow$ 0.79 $\pm$ 0.30 0.83 $\pm$ 0.27 0.83 $\pm$ 0.28 0.84 $\pm$ 0.27
Arc F1 $\uparrow$ 0.63 $\pm$ 0.39 0.70 $\pm$ 0.38 0.67 $\pm$ 0.39 0.71 $\pm$ 0.37
Circle F1 $\uparrow$ 0.84 $\pm$ 0.29 0.90 $\pm$ 0.23 0.88 $\pm$ 0.23 0.90 $\pm$ 0.22
Extrusion F1 $\uparrow$ 0.97 $\pm$ 0.08 0.98 $\pm$ 0.08 0.98 $\pm$ 0.07 0.98 $\pm$ 0.06
CD median $\downarrow$ 110.18 76.34 92.07 76.05
SIR $\downarrow$ 0.05 $\pm$ 0.16 0.05 $\pm$ 0.18 0.05 $\pm$ 0.18 0.04 $\pm$ 0.15
DangEL $\downarrow$ 1.36 $\pm$ 5.78 1.41 $\pm$ 6.05 1.54 $\pm$ 6.36 1.26 $\pm$ 6.55
SegE $\downarrow$ 0.59 $\pm$ 2.27 0.60 $\pm$ 2.46 0.62 $\pm$ 2.26 0.65 $\pm$ 3.85
FluxEE (×$\times 10^{2}$) $\downarrow$ 0.28 $\pm$ 3.46 0.31 $\pm$ 2.75 0.30 $\pm$ 3.08 0.38 $\pm$ 8.53
Watertightness (%) $\uparrow$ 86.62 86.50 85.21 87.35
@@
EECM $\uparrow$ 0.77 0.81 0.78 0.80
DMCD (×$\times 10^{3}$) $\downarrow$ 5.19 $\pm$ 10.57 4.09 $\pm$ 9.24 4.90 $\pm$ 10.69 4.43 $\pm$ 10.16
SD mean (×$\times 10^{2}$) $\downarrow$ 4.53 $\pm$ 10.18 3.18 $\pm$ 8.30 3.87 $\pm$ 9.31 3.45 $\pm$ 8.63
SD median (×$\times 10^{2}$) $\downarrow$ 0.00 0.00 0.00 0.00
@@ 4.3 Impact of Annotation Quality and Training/Evaluation Congruence
Conversely, both models exhibit a performance degradation when evaluated out-of-distribution. The model trained on CADmium annotations sees a notable drop in F1 scores and an increase in CD when tested on the Text2CAD dataset. Similarly, the model trained on Text2CAD annotations shows considerably worse geometric fidelity (e.g., higher CD) and higher IRR when evaluated on the CADmium dataset’s more natural prompts.
Table 4: Impact of annotation quality using the Text2CAD architecture. Models were trained on either CADmium or Text2CAD annotations and evaluated on both the CADmium and Text2CAD test sets. Metrics are calculated on the set of valid meshes generated by each model-dataset configuration individually.
Trained on CADmium Text2CAD
Evaluated on CADmium Text2CAD CADmium Text2CAD
@@
Line F1 $\uparrow$ 0.88 $\pm$ 0.22 0.73 $\pm$ 0.30 0.64 $\pm$ 0.25 0.82 $\pm$ 0.26
Arc F1 $\uparrow$ 0.58 $\pm$ 0.41 0.21 $\pm$ 0.36 0.24 $\pm$ 0.34 0.38 $\pm$ 0.43
Circle F1 $\uparrow$ 0.87 $\pm$ 0.27 0.57 $\pm$ 0.44 0.47 $\pm$ 0.39 0.76 $\pm$ 0.37
Extrusion F1 $\uparrow$ 0.98 $\pm$ 0.07 0.89 $\pm$ 0.18 0.81 $\pm$ 0.17 0.93 $\pm$ 0.15
CD median $\downarrow$ 0.33 51.67 34.23 0.38
SIR $\downarrow$ 0.05 $\pm$ 0.17 0.05 $\pm$ 0.17 0.08 $\pm$ 0.22 0.05 $\pm$ 0.17
DangEL $\downarrow$ 1.66 $\pm$ 7.13 1.16 $\pm$ 6.43 2.56 $\pm$ 9.55 1.13 $\pm$ 5.77
SegE $\downarrow$ 0.80 $\pm$ 6.09 0.53 $\pm$ 2.93 1.17 $\pm$ 4.53 0.55 $\pm$ 2.88
FluxEE (×$\times 10^{2}$) $\downarrow$ 0.46 $\pm$ 3.70 0.59 $\pm$ 5.84 0.97 $\pm$ 5.88 0.46 $\pm$ 4.61
Watertightness (%) $\uparrow$ 87.63 91.90 82.48 90.52
@@
EECM $\uparrow$ 0.86 0.76 0.66 0.82
DMCD (×$\times 10^{3}$) $\downarrow$ 3.65 $\pm$ 9.00 8.29 $\pm$ 13.16 9.41 $\pm$ 12.10 5.35 $\pm$ 10.77
SD mean (×$\times 10^{2}$) $\downarrow$ 2.71 $\pm$ 6.88 8.69 $\pm$ 12.19 11.52 $\pm$ 15.24 5.04 $\pm$ 10.32
SD median (×$\times 10^{2}$) $\downarrow$ 0.00 3.65 5.33 0.11

### v2

@@ 4 Experiments
This section provides an overview of our experimental setup and presents the evaluation of the proposed CAD generation pipeline. We conduct experiments across four datasets: the original Text2CAD dataset (Khan et al., 2024), our re-annotated DeepCAD dataset (CADmium), the CADPrompt dataset (Alrashedy et al., 2024) containing 200 expert-annotated objects from DeepCAD (Wu et al., 2021), and the Fusion360 Reconstruction dataset (Willis et al., 2021) annotated using our pipeline.
@@
Additionally, for the Chamfer Distance, we observed that rare outliers with extremely high values skewed the mean, rendering it less informative. We therefore report the median CD as a more robust alternative. All metrics are accompanied by their $95\%$ confidence interval computed over individual test samples for a single model run.
To assess whether observed differences between models are statistically significant, we apply different hypothesis tests depending on the evaluation protocol. When comparing results on non-identical sets of valid meshes (main paper tables), we treat the samples as independent and use unpaired tests: Welch’s t-test if both groups appear approximately normal, otherwise the Mann–Whitney U test, and for binary outcomes such as validity and watertightness we use Fisher’s exact test or the Chi-square test with Yates’ correction when expected counts are sufficient. When comparing results on the common subset of objects valid across models (Appendix B), we use paired tests: the Wilcoxon signed-rank test for continuous metrics and McNemar’s exact test for binary outcomes. In all cases, we report differences as statistically significant if the adjusted p-value (Benjamini–Hochberg correction within each metric) is below 0.05.
@@ 4.1 Generalization to Human-Annotated Prompts (CADPrompt)
To assess our models’ ability to generalize to prompts authored by humans, we evaluated them on the CADPrompt dataset (Alrashedy et al., 2024). This dataset comprises expert-written natural language instructions for 200 DeepCAD samples, serving as a crucial test for out-of-distribution generalization. Table 1 presents the results, comparing our Qwen Coder 14B model trained on CADmium data against the original Text2CAD model as a key baseline. A corresponding common subset analysis is provided in Appendix B.
@@
Table 1: Performance comparison on the human-annotated CADPrompt dataset. Metrics are calculated on the set of valid meshes generated by each model individually. Reported values include 95% confidence intervals in brackets. Boldface indicates the better-performing model when the difference is statistically significant at $p<0.05$. Boldface indicates statistically significant differences ($p<0.05$); details of the statistical testing procedure are provided in Section 4.
Model Text2CAD Qwen-2.5 Coder 14B
@@
Line F1 $\uparrow$ 0.67 [$0.62$, $0.71$] 0.69 [$0.64$, $0.74$]
Arc F1 $\uparrow$ 0.00 [$0.00$, $0.00$] 0.22 [$0.11$, $0.35$]
Circle F1 $\uparrow$ 0.31 [$0.23$, $0.41$] 0.61 [$0.51$, $0.71$]
Extrusion F1 $\uparrow$ 0.84 [$0.82$, $0.87$] 0.89 [$0.87$, $0.92$]
CD mean $\downarrow$ 221.42 [$185.79$, $258.88$] 205.46 [$175.29$, $237.00$]
CD median $\downarrow$ 127.70 [$87.59$, $162.91$] 116.75 [$81.64$, $179.09$]
SIR $\downarrow$ 0.02 [$0.01$, $0.04$] 0.11 [$0.07$, $0.15$]
DangEL $\downarrow$ 0.20 [$0.03$, $0.44$] 2.98 [$1.67$, $4.55$]
SegE $\downarrow$ 0.09 [$0.03$, $0.17$] 0.78 [$0.48$, $1.14$]
FluxEE ($\times 10^{2}$) $\downarrow$ 0.04 [$0.00$, $0.12$] 2.24 [$1.01$, $3.73$]
Watertightness (%) $\uparrow$ 95.74 [$92.55$, $98.40$] 83.80 [$78.21$, $88.83$]
@@
EECM $\uparrow$ 0.66 [$0.59$, $0.73$] 0.68 [$0.60$, $0.75$]
DMCD ($\times 10^{3}$) $\downarrow$ 13.39 [$11.61$, $15.26$] 8.42 [$6.63$, $10.35$]
SD mean ($\times 10^{2}$) $\downarrow$ 14.78 [$12.74$, $17.02$] 7.02 [$5.05$, $9.26$]
SD median ($\times 10^{2}$) $\downarrow$ 11.15 [$8.68$, $13.64$] 1.54 [$1.07$, $2.15$]
@@ 4.2 Impact of Model Scale
A primary benefit of increasing model size is the consistent reduction in the Invalidity Ratio (IR). On the CADmium test set, IR drops progressively with larger models, and the 14B model achieves a notably lower IR on the more challenging Fusion360 dataset as well. Accuracy in reconstructing geometric primitives, as measured by F1 scores, generally improves or remains robust with increasing model size, particularly when compared on a common subset of generated objects (see Appendix B). Metrics related to fine-grained mesh quality, such as DangEL and SD, also tend to show improvements with larger models, suggesting better adherence to manifold properties. Some examples of the reconstructions generated by the 14B model are shown in Figure˜3.
Table 2: Impact of Qwen Coder model scale (1.5B, 3B, 7B, 14B parameters), trained on CADmium annotations, evaluated on the CADmium test set. Metrics are calculated on the set of valid meshes generated by each model individually for each dataset. Reported values include 95% confidence intervals in brackets. Statistical significance is assessed with statistical tests on adjacent sizes (1.5B→3B, 3B→7B, 7B→14B). Bold indicates a size that is significantly better than the next smaller size at $p<0.05$. Details of the statistical testing procedure are provided in Section 4.
@@
Line F1 $\uparrow$ 0.90 [$0.90$, $0.91$] 0.91 [$0.90$, $0.91$] 0.91 [$0.90$, $0.91$] 0.92 [$0.91$, $0.92$]
Arc F1 $\uparrow$ 0.70 [$0.68$, $0.73$] 0.74 [$0.72$, $0.76$] 0.73 [$0.71$, $0.75$] 0.78 [$0.76$, $0.80$]
Circle F1 $\uparrow$ 0.87 [$0.86$, $0.88$] 0.91 [$0.90$, $0.92$] 0.89 [$0.89$, $0.90$] 0.92 [$0.91$, $0.92$]
Extrusion F1 $\uparrow$ 0.98 [$0.98$, $0.98$] 0.98 [$0.98$, $0.98$] 0.98 [$0.98$, $0.99$] 0.99 [$0.98$, $0.99$]
CD mean $\downarrow$ 70.48 [$66.65$, $74.52$] 43.91 [$40.52$, $47.36$] 63.11 [$59.41$, $66.91$] 42.69 [$39.71$, $45.85$]
CD median $\downarrow$ 0.31 [$0.29$, $0.32$] 0.24 [$0.23$, $0.25$] 0.27 [$0.26$, $0.28$] 0.23 [$0.22$, $0.25$]
SIR $\downarrow$ 0.04 [$0.04$, $0.04$] 0.05 [$0.04$, $0.05$] 0.04 [$0.04$, $0.05$] 0.04 [$0.04$, $0.04$]
DangEL $\downarrow$ 1.18 [$1.03$, $1.35$] 1.35 [$1.20$, $1.50$] 1.10 [$0.98$, $1.23$] 1.13 [$1.00$, $1.27$]
SegE $\downarrow$ 0.48 [$0.42$, $0.55$] 0.52 [$0.46$, $0.57$] 0.48 [$0.43$, $0.53$] 0.54 [$0.45$, $0.67$]
FluxEE ($\times 10^{2}$) $\downarrow$ 0.38 [$0.30$, $0.47$] 0.49 [$0.39$, $0.60$] 0.36 [$0.29$, $0.43$] 0.39 [$0.31$, $0.48$]
Watertightness (%) $\uparrow$ 90.38 [$89.71$, $91.05$] 90.31 [$89.63$, $90.99$] 90.54 [$89.88$, $91.20$] 90.78 [$90.13$, $91.42$]
@@
EECM $\uparrow$ 0.87 [$0.86$, $0.88$] 0.89 [$0.88$, $0.90$] 0.88 [$0.87$, $0.89$] 0.89 [$0.88$, $0.90$]
DMCD ($\times 10^{3}$) $\downarrow$ 2.97 [$2.78$, $3.16$] 2.68 [$2.49$, $2.87$] 2.73 [$2.54$, $2.93$] 2.67 [$2.48$, $2.86$]
SD mean ($\times 10^{2}$) $\downarrow$ 2.90 [$2.70$, $3.11$] 2.19 [$2.02$, $2.37$] 2.23 [$2.06$, $2.41$] 1.99 [$1.83$, $2.15$]
SD median ($\times 10^{2}$) $\downarrow$ 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$]
Table 3: Impact of Qwen Coder model scale (1.5B, 3B, 7B, 14B parameters), trained on CADmium annotations, evaluated on the Fusion360 test set. Metrics are calculated on the set of valid meshes generated by each model individually for each dataset. Reported values include 95% confidence intervals in brackets. Statistical significance is assessed with statistical tests on adjacent sizes (1.5B→3B, 3B→7B, 7B→14B). Bold indicates a size that is significantly better than the next smaller size at $p<0.05$. Details of the statistical testing procedure are provided in Section 4.
@@
Line F1 $\uparrow$ 0.79 [$0.78$, $0.80$] 0.83 [$0.83$, $0.84$] 0.83 [$0.82$, $0.83$] 0.84 [$0.83$, $0.85$]
Arc F1 $\uparrow$ 0.63 [$0.62$, $0.65$] 0.70 [$0.68$, $0.71$] 0.67 [$0.65$, $0.68$] 0.71 [$0.70$, $0.73$]
Circle F1 $\uparrow$ 0.84 [$0.83$, $0.85$] 0.90 [$0.89$, $0.90$] 0.88 [$0.87$, $0.88$] 0.90 [$0.89$, $0.90$]
Extrusion F1 $\uparrow$ 0.97 [$0.97$, $0.98$] 0.98 [$0.97$, $0.98$] 0.98 [$0.98$, $0.98$] 0.98 [$0.98$, $0.98$]
CD mean $\downarrow$ 226.46 [$218.86$, $234.38$] 191.94 [$184.22$, $199.68$] 206.39 [$199.19$, $213.78$] 202.00 [$194.52$, $209.37$]
CD median $\downarrow$ 110.18 [$104.45$, $117.15$] 76.34 [$69.24$, $83.42$] 92.07 [$86.51$, $97.20$] 76.05 [$70.08$, $82.17$]
SIR $\downarrow$ 0.05 [$0.04$, $0.05$] 0.05 [$0.05$, $0.06$] 0.05 [$0.05$, $0.06$] 0.04 [$0.04$, $0.05$]
DangEL $\downarrow$ 1.36 [$1.23$, $1.51$] 1.41 [$1.27$, $1.56$] 1.54 [$1.40$, $1.70$] 1.26 [$1.12$, $1.42$]
SegE $\downarrow$ 0.59 [$0.54$, $0.65$] 0.60 [$0.55$, $0.67$] 0.62 [$0.56$, $0.67$] 0.65 [$0.57$, $0.75$]
FluxEE ($\times 10^{2}$) $\downarrow$ 0.28 [$0.21$, $0.38$] 0.31 [$0.25$, $0.38$] 0.30 [$0.23$, $0.38$] 0.38 [$0.24$, $0.61$]
Watertightness (%) $\uparrow$ 86.62 [$85.80$, $87.48$] 86.50 [$85.67$, $87.35$] 85.21 [$84.38$, $86.05$] 87.35 [$86.57$, $88.09$]
@@
EECM $\uparrow$ 0.77 [$0.76$, $0.78$] 0.81 [$0.80$, $0.82$] 0.78 [$0.77$, $0.79$] 0.80 [$0.79$, $0.81$]
DMCD ($\times 10^{3}$) $\downarrow$ 5.19 [$4.91$, $5.48$] 4.09 [$3.84$, $4.35$] 4.90 [$4.62$, $5.19$] 4.43 [$4.18$, $4.69$]
SD mean ($\times 10^{2}$) $\downarrow$ 4.53 [$4.25$, $4.81$] 3.18 [$2.96$, $3.41$] 3.87 [$3.63$, $4.12$] 3.45 [$3.23$, $3.68$]
SD median ($\times 10^{2}$) $\downarrow$ 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$]
@@ 4.3 Impact of Annotation Quality and Training/Evaluation Congruence
Conversely, both models exhibit a performance degradation when evaluated out-of-distribution. The model trained on CADmium annotations sees a notable drop in F1 scores and an increase in CD when tested on the Text2CAD dataset. Similarly, the model trained on Text2CAD annotations shows considerably worse geometric fidelity (e.g., higher CD) and higher IR when evaluated on the CADmium dataset’s more natural prompts.
Table 4: Impact of annotation quality using the Text2CAD architecture. Models were trained on either CADmium or Text2CAD annotations and evaluated on both the CADmium and Text2CAD test sets. Metrics are calculated on the set of valid meshes generated by each model–dataset configuration individually. Reported values include 95% confidence intervals in brackets. Boldface indicates statistically significant differences ($p<0.05$); details of the statistical testing procedure are provided in Section 4.
Evaluated on CADmium Text2CAD
Trained on CADmium Text2CAD CADmium Text2CAD
@@
Line F1 $\uparrow$ 0.88 [$0.87$, $0.88$] 0.73 [$0.72$, $0.74$] 0.64 [$0.63$, $0.64$] 0.82 [$0.81$, $0.83$]
Arc F1 $\uparrow$ 0.58 [$0.56$, $0.60$] 0.21 [$0.19$, $0.23$] 0.24 [$0.23$, $0.26$] 0.38 [$0.36$, $0.40$]
Circle F1 $\uparrow$ 0.87 [$0.86$, $0.88$] 0.57 [$0.56$, $0.59$] 0.47 [$0.46$, $0.48$] 0.76 [$0.74$, $0.77$]
Extrusion F1 $\uparrow$ 0.98 [$0.98$, $0.98$] 0.89 [$0.89$, $0.90$] 0.81 [$0.80$, $0.81$] 0.93 [$0.93$, $0.94$]
CD mean $\downarrow$ 53.55 [$50.17$, $57.06$] 145.87 [$141.11$, $150.64$] 107.99 [$104.16$, $111.93$] 29.38 [$27.50$, $31.35$]
CD median $\downarrow$ 0.33 [$0.32$, $0.36$] 51.67 [$47.72$, $56.13$] 34.23 [$31.87$, $36.29$] 0.38 [$0.35$, $0.42$]
SIR $\downarrow$ 0.05 [$0.05$, $0.05$] 0.05 [$0.04$, $0.05$] 0.08 [$0.08$, $0.09$] 0.05 [$0.04$, $0.05$]
DangEL $\downarrow$ 1.66 [$1.50$, $1.83$] 1.16 [$1.02$, $1.31$] 2.56 [$2.34$, $2.78$] 1.13 [$1.01$, $1.26$]
SegE $\downarrow$ 0.80 [$0.68$, $0.95$] 0.53 [$0.47$, $0.61$] 1.17 [$1.08$, $1.28$] 0.55 [$0.49$, $0.62$]
FluxEE ($\times 10^{2}$) $\downarrow$ 0.46 [$0.39$, $0.55$] 0.59 [$0.46$, $0.73$] 0.97 [$0.84$, $1.10$] 0.46 [$0.36$, $0.57$]
Watertightness (%) $\uparrow$ 87.63 [$86.90$, $88.36$] 91.90 [$91.29$, $92.50$] 82.48 [$81.63$, $83.33$] 90.52 [$89.85$, $91.17$]
@@
EECM $\uparrow$ 0.86 [$0.85$, $0.87$] 0.76 [$0.75$, $0.77$] 0.66 [$0.65$, $0.67$] 0.82 [$0.81$, $0.83$]
DMCD ($\times 10^{3}$) $\downarrow$ 3.65 [$3.44$, $3.87$] 8.29 [$7.98$, $8.62$] 9.41 [$9.11$, $9.73$] 5.35 [$5.10$, $5.61$]
SD mean ($\times 10^{2}$) $\downarrow$ 2.71 [$2.54$, $2.87$] 8.69 [$8.39$, $8.99$] 11.52 [$11.14$, $11.91$] 5.04 [$4.79$, $5.29$]
SD median ($\times 10^{2}$) $\downarrow$ 0.00 [$0.00$, $0.00$] 3.65 [$3.41$, $4.01$] 5.33 [$4.90$, $5.69$] 0.11 [$0.06$, $0.16$]